A plot of land,600sq m in area ,is divided between two persons such that the first persons gets three-fifth of what the second gets. Find the share of each.

Answer by Boreal(15235)

600 m^2
x is what one person gets
second gets 5/3 x
(3/5)(5/3)=1
x+(5/3)x=600
(8/3)x=600
x=600*(3/8)=225
(5/3)x=375
(225/375)=9/15=3/5

225m^2; 375 m^2